Find the cost per pound of a coffee mixtur made from 7 lbs of coffee that costs $9.80 per lb and 3lbs of coffee that costa $8.10 per lb?
Answer by fractalier(2101)

We find the average this way
[7*$9.80 + 3*$8.10] / 10 =
(68.60 + 24.30) / 10 =
$92.90 / 10 =
$9.29 per pound